### v3

# Convergence Rate of the Join-the-Shortest-Queue System

[…]
Convergence Rate of the Join-the-Shortest-Queue System
[…]
The Join-the-Shortest-Queue (JSQ) policy is among the most widely used load balancing algorithms and has been extensively studied. However, an exact characterization of the system behavior remains challenging. Most prior research has focused on analyzing its performance in the steady state in certain asymptotic regimes, such as the heavy-traffic regime. However, the convergence rate to the steady state in these regimes is often slow, so steady-state and heavy-traffic characterizations may be less informative over practical time horizons. To address this limitation, we provide a finite-time convergence rate analysis of a JSQ system with two symmetric servers. In sharp contrast to the existing literature, we directly study the original system rather than an approximate limiting system such as a diffusion approximation. Our results demonstrate that for such a system, the convergence rate to its steady state, measured in the total variation distance, is $O\left(\frac{1}{(1-\rho)^{3}}\frac{1}{t}\right)$, where $\rho\in(0,1)$ is the traffic intensity.
[…]
Join-the-Shortest-Queue; Load Balancing System; Transient Analysis; Coupling
1 Introduction
Efficient load balancing is central to optimizing resource allocation and minimizing job delays in modern service systems and data centers. The Join–the-Shortest-Queue (JSQ) policy is a popular load balancing algorithm due to its simplicity and optimality under various criteria. Since its introduction by Winston [20], it has received significant attention over the decades. However, the analysis of the JSQ system remains highly nontrivial [5]. Due to the complexity of the problem, most prior works consider a JSQ system only in steady state. Further, researchers typically conduct analysis in some asymptotic regimes, such as the heavy-traffic regime. For example, there are various bounds for the queue length of a JSQ system in its steady state, such as [15, Theorem 10.2.3]. However, traffic patterns in modern data centers tend to be time-varying [2]. This necessitates careful finite-time analysis of JSQ performance, rather than relying only on steady-state behavior. Motivated by this, we focus on analyzing the transient behavior of a JSQ system. As a starting point, we consider a system consisting of two servers with both interarrival and service times following exponential distributions. We analyze its non-asymptotic convergence rate toward its steady state.
As discussed, prior works focus on analyzing the JSQ policy under the heavy-traffic regime and in its steady state. However, unsurprisingly, simulation results in Figure 1 reveal that the convergence rate of a JSQ system can be quite slow when the traffic intensity $\rho$ approaches 1. Thus, to conduct more careful analysis of the JSQ systems, we need to understand their transient behavior. In addition, prior results are not applicable to light-traffic settings since they rely on state space collapse (SSC), which is only associated with the heavy-traffic regime. In this paper, we address this challenge and obtain an explicit upper bound on the difference between the transient distribution of a JSQ system and its steady state for any traffic intensity $\rho\in(0,1)$.
Figure 1: Finite-time performance of a JSQ system at different values of $\rho$, where $q_{i}(t)$ denotes the queue length at time $t$ for server ii. The results are based on 1000 simulations of trajectories to estimate the expected total queue length. We consider a JSQ system with a Poisson arrival process of rate $\lambda=\rho$ and two identical servers, each with exponentially distributed service times of rate $\mu=0.5$. Results are shown for $\rho=0.5$ (left) and $\rho=0.999$ (right). The pair of plots clearly shows that the convergence rates of JSQ systems heavily depend on $\rho$.
We focus on analyzing the distribution $\mathbb{P}^{t}_{\bm{x}}$ of the queue length vector at time $t$, for a JSQ system with initial queue length $\bm{x}$. In particular, we consider its total variation (TV) distance to the steady state $\pi$ at every time $t\geq 0$. The total variation distance is a common metric quantifying the convergence rate of a stochastic system and is known to be closely connected to many other metrics. For example, bounds on TV distances imply bounds on Hellinger distances. In addition to the TV distance, we derive the convergence rate of the mean queue length. Using coupling, the key challenge in our method is to obtain an upper bound on the expected hitting times or first passage times of the JSQ system. For general continuous-time Markov chains (CTMCs), analyzing the hitting time usually requires extensive exploitation of the structure of the CTMC [4]. We use several structures of the JSQ system, and show that the problem can be reduced to providing an upper bound on $\mathbb{E}[T_{0}]$, where $T_{0}$, to be defined in (3.1), is the time it takes for a JSQ system’s queue length vector to hit the origin starting from state $(1,0)$. After upper bounding $\mathbb{E}[T_{0}]$, we use Markov’s inequality to arrive at a bound on $d_{\mathrm{TV}}\left(\mathbb{P}^{t}_{\bm{x}},\pi\right)$.
1.1 Main contribution
Our main result is for a continuous-time JSQ system with two symmetric queues, where both interarrival and service times follow exponential distributions. We derive an upper bound on $d_{\mathrm{TV}}\left(\mathbb{P}^{t}_{\bm{x}},\pi\right)$ and, as a corollary, an upper bound on the difference between the mean queue length at time $t$ and its steady-state value. These two distances converge to zero as $t\to\infty$ for any traffic intensity $\rho\in(0,1)$. In particular, for the TV distance, we establish in Theorem 1 that the convergence rate is $O\left(\frac{1}{(1-\rho)^{3}}\frac{1}{t}\right).$ In contrast to prior work based on diffusion limits, we derive transient bounds by directly studying the original JSQ system, thus avoiding approximation errors. To the best of our knowledge, we are the first to provide a convergence rate analysis of a JSQ system for any $\rho\in(0,1)$.
The rest of the paper is organized as follows. We first discuss related work in Section 1.2 and notation in Section 1.3. We present the model and state our main result in Section 2. We provide all proofs in Section 3. Finally, we provide concluding remarks and promising future directions in Section 4.
1.2 Related work
Since the pioneering work of Winston [20], the JSQ policy has been extensively studied in the literature [3, 5]. Transient analyses are generally complicated, even for simple systems such as $M/M/1$ queues [16, 1] and $M/M/n$ queues [8]. There are numerical results for transient analysis of JSQ systems [9], but few theoretical results are known regarding their convergence rates. In this work, we analyze the transient behavior through a hitting time analysis (see, e.g., [10]). Though there are some results related to hitting times or first passage times for JSQ systems [21, 17], they cannot directly be used for our problem. For a simple system such as an $M/M/1$ system, one can obtain the moment generating function (MGF) of the hitting time for the queue length to hit state 0 using a construction of an exponential martingale and the optional stopping theorem [14, Proposition 5.4]. One of the key reasons this works is the special property of an $M/M/1$ queue: starting from state $x\geq 1$, the queue length must first hit states $x-1,x-2,\cdots,1$ before hitting state 0. However, this property is not true for a JSQ system. For example, consider a JSQ system with an initial queue length $(x,y)$ with $x,y\geq 1$. It is possible that the queue length vector never hits state (x1,y1)(x-1,y-1) before hitting $(0,0)$. This is one of the technical barriers to analyzing a JSQ system. One possible approach is to use the exponential martingale constructed in [17]; however, that work focuses on the hitting time to $(K,K)$, $K\geq 1$, whereas our interest is in the hitting time to $(0,0)$, which is fundamentally different, making those results inapplicable.
The convergence rates of stochastic systems and hitting time analyses based on Foster–Lyapunov drift functions have been extensively studied in the literature; see, e.g., [7]. For example, Theorem 2 of [12] provides a general framework for obtaining exponential convergence rates of general stochastic systems. However, applying it requires a suitable drift function, and it is unclear how to construct one for the JSQ system. Relatedly, researchers have discussed transient analysis for queues under various limiting regimes or for simpler systems [19, 8]. Compared to these works, we analyze the JSQ system in a fixed system setting.
1.3 Notation
In this section, we introduce the notation that will be used throughout this paper. We use $\mathbb{P}\left(A\right)$ to denote the probability that event $A$ occurs, and use $\mathbb{E}[X]$ to denote the expected value of the random variable $X$. For a vector $\bm{x}$, we use $x_{i}$ to denote its $i$-th element. Given two vectors $\bm{x}$ and $\bm{y}$ in $\mathbb{R}^{n}$, we write $\bm{x}\geq\bm{y}$ if $x_{i}\geq y_{i}$ for every $i\in\{1,\ldots,n\}$. For a stochastic process $(X_{t})_{t\geq 0}$ with steady state $\pi$, we use $\mathbb{E}_{\pi}[X]$ and $\mathbb{P}_{\pi}(\cdot)$ to denote the expectation and distribution of $X_{t}$ with respect to the steady state. For two random variables $A$ and $B$, we write that $A$ and $B$ are equal in distribution, or $A{\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}B$, if $\mathbb{P}\left(A\leq x\right)=\mathbb{P}\left(B\leq x\right)$ for all $x\in\mathbb{R}$. We define $A$ to be stochastically dominated by $B$, or $A\preceq B$, if and only if $\mathbb{P}\left(A>x\right)\leq\mathbb{P}\left(B>x\right)$ for all $x\in\mathbb{R}$. For two probability distributions $\mathbb{P}$ and $\mathbb{Q}$ defined on the sigma-algebra $\mathcal{F}$, we denote the total variation distance between $\mathbb{P}$ and $\mathbb{Q}$ as $d_{\mathrm{TV}}\left(\mathbb{P},\mathbb{Q}\right)=\sup_{A\in\mathcal{F}}|\mathbb{P}(A)-\mathbb{Q}(A)|.$
2 Model and results
2.1 Model
[…]
where $\bm{e}^{(i)}$ is the $i$-th unit vector.
2.2 Main result
We now present the main result of this paper: the convergence rate for the JSQ system described in Section 2.1.
Theorem 1
Consider a JSQ system as described in Section 2.1 with $\rho=\frac{\lambda}{2\mu}<1$. Let its initial queue length vector be $\bm{x}$, and let $\mathbb{P}_{\bm{x}}^{t}$ be the distribution of its queue length vector $\bm{q}(t)$ at time $t$, and let $\pi$ be its steady state. Then, for all $t>0$,
$d_{\mathrm{TV}}\left(\mathbb{P}_{\bm{x}}^{t},\pi\right)\leq\frac{1}{t}C_{1}(\lambda,\mu,\bm{x}),$ (2.1)
where
$C_{1}(\lambda,\mu,\bm{x})=\frac{2-\rho}{2\mu(1-\rho)^{3}}\Big((x_{1}+x_{2})(1-\rho)+\rho(2-\rho)\Big).$ (2.2)
Further, define
$K(\rho)=\begin{cases}\min\left\{1,\frac{2+2\rho}{1-2\rho^{2}}\cdot\frac{(1-\rho)^{2}}{2-\rho}\right\},&\mathrm{\penalty 10000\ if\penalty 10000\ }\rho\in[0,\frac{1}{\sqrt{2}}),\\ 1,&\mathrm{\penalty 10000\ if\penalty 10000\ }\rho\in[\frac{1}{\sqrt{2}},1),\end{cases}$ (2.3)
and let
$C_{1}^{K}(\lambda,\mu,\bm{x})=C_{1}(\lambda,\mu,\bm{x})\cdot K(\rho).$ (2.4)
Then, for all $t>0$,
$d_{\mathrm{TV}}\left(\mathbb{P}_{\bm{x}}^{t},\pi\right)\leq\frac{1}{t}C_{1}^{K}(\lambda,\mu,\bm{x}).$ (2.5)
Figure 2: Left: plot of the bound (2.5) as a function of $t$ for different values of $\rho$ and $x_{1}=x_{2}=10,\mu=1$. Right: plot of $K(\rho)$ (2.3) as a function of $\rho$.
We now visualize the upper bound (2.5) and the value of $K(\rho)$ (2.3) in Figure 2. As we observe, the right-hand side of (2.5) varies significantly across $\rho$, though its dependence on $t$ is the same for all $\rho$.
Next, in Corollary 1, we establish a convergence rate result for the difference between the transient and steady-state mean queue lengths. In particular, the bound (2.6) implies that the mean queue length converges at rate $O\left(\frac{1}{(1-\rho)^{3}}\frac{1}{\sqrt{t}}\right)$; the proof is provided in Section 3.9. Combining this corollary with existing results on the stationary queue length of the JSQ system, such as Proposition 4 of [6], yields an explicit transient upper bound on the mean queue length, as stated in (2.7).
Corollary 1
Consider a JSQ system as described in Theorem 1 with $\rho=\frac{\lambda}{2\mu}<1$ and initial queue length vector $\bm{x}$. Then
$|\mathbb{E}[q_{1}(t)+q_{2}(t)]-\mathbb{E}_{\pi}[q_{1}+q_{2}]|\leq\frac{1}{\sqrt{t}}\Big(2C_{1}^{K}(\lambda,\mu,\bm{x})+C_{2}(\lambda,\mu,\bm{x})\Big)+\frac{2}{t}C_{1}^{K}(\lambda,\mu,\bm{x})$ (2.6)
and
$\mathbb{E}[q_{1}(t)+q_{2}(t)]\leq\frac{\rho(2-\rho)}{1-\rho}+\frac{1}{\sqrt{t}}\Big(2C_{1}^{K}(\lambda,\mu,\bm{x})+C_{2}(\lambda,\mu,\bm{x})\Big)+\frac{2}{t}C_{1}^{K}(\lambda,\mu,\bm{x}),$ (2.7)
where $C_{1}^{K}(\lambda,\mu,\bm{x})$ is defined in (2.4) and
$C_{2}(\lambda,\mu,\bm{x})=(x_{1}^{2}+x_{2}^{2})+2(x_{1}+x_{2})\frac{\rho}{1-\rho}+\frac{4\rho(1+\rho)}{(1-\rho)^{2}}.$ (2.8)
3 Proof of Theorem 1
In the proof, we use the following hitting time, denoted by $T_{0}$ (3.1). Consider a JSQ system described in Section 2.1 with queue length process $\bm{q}(t)$. We define
$T_{0}$ $=\inf\left\{t\geq 0:\bm{q}(t)=(0,0)\mid\bm{q}(0)=(1,0)\right\}$ (3.1)
as the time for the queue length process to hit state $(0,0)$ starting from $(1,0)$. In the proof, we use two coupled processes defined in Section 3.1.
3.1 Coupling with arrival and service
Consider two JSQ systems with queue length vectors $\bm{p}(t)$ and $\bm{q}(t)$ as described in Section 2.1. We call them coupled with arrival and service if they can be constructed using $\mathcal{N}_{a}$ and $\mathcal{N}_{d}^{i}$, $i=1,2$, where $\mathcal{N}_{a}$ is a Poisson process with rate $\lambda$, and $\mathcal{N}_{d}^{1},\mathcal{N}_{d}^{2}$ are two independent Poisson processes, each with rate $\mu$. We allow the initial queue length vectors of $\bm{p}$ and $\bm{q}$ to be different. We also assume that $\mathcal{N}_{a}$, $\mathcal{N}_{d}^{1}$, and $\mathcal{N}_{d}^{2}$ are independent of each other. We interpret $\mathcal{N}_{a}$ as the arrival process and $\mathcal{N}_{d}^{1},\mathcal{N}_{d}^{2}$ as the potential service processes. Accordingly, we can write the expressions for the dynamics of the queue length vectors in the two JSQ systems as
$\begin{split}dp_{1}(t)=\mathbb{I}_{\left\{p_{1}(t-)\leq p_{2}(t-)\right\}}\mathcal{N}_{a}(dt)-\mathbb{I}_{\left\{p_{1}(t-)\geq 1\right\}}\mathcal{N}_{d}^{1}(dt),\\ dp_{2}(t)=\mathbb{I}_{\left\{p_{1}(t-)>p_{2}(t-)\right\}}\mathcal{N}_{a}(dt)-\mathbb{I}_{\left\{p_{2}(t-)\geq 1\right\}}\mathcal{N}_{d}^{2}(dt),\\ dq_{1}(t)=\mathbb{I}_{\left\{q_{1}(t-)\leq q_{2}(t-)\right\}}\mathcal{N}_{a}(dt)-\mathbb{I}_{\left\{q_{1}(t-)\geq 1\right\}}\mathcal{N}_{d}^{1}(dt),\\ dq_{2}(t)=\mathbb{I}_{\left\{q_{1}(t-)>q_{2}(t-)\right\}}\mathcal{N}_{a}(dt)-\mathbb{I}_{\left\{q_{2}(t-)\geq 1\right\}}\mathcal{N}_{d}^{2}(dt),\end{split}$ (3.2)
where $p_{1}(t-)$ is the value of $p_{1}$ evaluated just before time $t$.
3.2 Partial proof of Theorem 1
We first prove a weaker version of Theorem 1: we only prove it up to (2.1). The proof builds on Proposition 1 and Lemma 1. We present the complete proof of Theorem 1 in Section 3.7.
Proposition 1
[10, 11] Let $(X_{t})_{t\geq 0}$ and $(Y_{t})_{t\geq 0}$ be two copies of the same continuous-time, discrete-valued Markov chain on a common probability space with $X_{0}=\bm{x}$ and $Y_{0}\sim\pi$, where $\pi$ is the steady state of this Markov chain. We assume $X_{t}$ and $Y_{t}$ take values in a countable set $\mathcal{X}$, and let $T=\inf\{t\geq 0:X_{t}=Y_{t}\}$ denote the first time that they meet. Then
$d_{\mathrm{TV}}\left(\mathbb{P}_{\bm{x}}^{t},\pi\right)\leq\mathbb{P}\left(T\geq t\right).$
Lemma 1
Consider two coupled JSQ systems with queue length processes $\bm{p}(t)$ and $\bm{q}(t)$ as governed by (3.2). Assume $\bm{p}(0)=\bm{x}$ and $\bm{q}(0){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}\pi$, where $\pi$ is the steady state of $\bm{p}$. Let $T$ be the first time that the two systems have the same queue length vector. We have
$\mathbb{P}\left(T\geq t\right)\leq\frac{1}{t}C_{1}(\lambda,\mu,\bm{x}),$ (3.3)
where $C_{1}(\lambda,\mu,\bm{x})$ is defined in (2.2).
Now we are ready to present the proof of Theorem 1 up to (2.1).
Proof of Theorem 1 up to (2.1).
Applying Proposition 1 yields
$d_{\mathrm{TV}}\left(\mathbb{P}_{\bm{x}}^{t},\pi\right)\leq\mathbb{P}\left(T\geq t\right),$
where $T$ is defined in Lemma 1. By (3.3) in Lemma 1,
$d_{\mathrm{TV}}\left(\mathbb{P}_{\bm{x}}^{t},\pi\right)\leq\mathbb{P}\left(T\geq t\right)\leq\frac{1}{t}C_{1}(\lambda,\mu,\bm{x})$
and we complete the proof of (2.1). ∎
3.3 Proof of Lemma 1
To prove Lemma 1, we first establish a monotonicity property for coupled JSQ systems in Lemma 2 with proof in Section 3.4. Next, we discuss Lemma 3, proved in Section 3.5, which provides an upper bound on the hitting time to the origin for a JSQ system starting from an arbitrary state $\bm{x}$, in terms of $T_{0}$ defined in (3.1). We then bound $\mathbb{E}[T_{0}]$ in Lemma 4 with proof given in Section 3.6. Together, these results imply Lemma 1.
Lemma 2
Consider two coupled JSQ systems (3.2) with queue length processes $\bm{p}(t)$ and $\bm{q}(t)$. If $p_{i}(0)\geq q_{i}(0)$ for $i\in\{1,2\}$, then $p_{i}(t)\geq q_{i}(t)$ for $i\in\{1,2\}$ and for every $t\geq 0$. In particular, for any $i\in\{1,2\}$ and $t\geq 0$, $p_{i}(t)=0\Rightarrow q_{i}(t)=0$. Further, define $r_{i}(t)=p_{i}(t)-q_{i}(t)\geq 0$ and $r(t)=r_{1}(t)+r_{2}(t)$. Then $r(t)\leq r(0)$ for any $t\geq 0.$
Lemma 3
Consider a JSQ system with queue length process $\bm{q}(t)$. Define the random variable $\mathcal{T}(\bm{x})$ as the time for $\bm{q}(t)$ to first hit state $(0,0)$ with $\bm{q}(0)=(x_{1},x_{2})$:
$\mathcal{T}(\bm{x})=\mathcal{T}(x_{1},x_{2})=\inf\{t\geq 0:\bm{q}(t)=(0,0)|\bm{q}(0)=(x_{1},x_{2})\}.$
Then we have $\mathcal{T}(x_{1},x_{2})\preceq\sum_{i=1}^{x_{1}+x_{2}}T_{0i}$, where $T_{0i}$ are independent copies of $T_{0}$ (3.1) and the relation $\preceq$ is defined in Section 1.3.
Lemma 4
Consider $T_{0}$ (3.1) and assume $\rho=\frac{\lambda}{2\mu}<1$. Then
$\mathbb{E}[T_{0}]\leq\frac{1}{\lambda}\frac{\rho(2-\rho)}{(1-\rho)^{2}}.$ (3.4)
To prove Lemma 1, we also use Proposition 2 from the literature, which provides an upper bound on $\mathbb{E}_{\pi}[q_{1}+q_{2}]$ for the JSQ system.
Proposition 2 (Proposition 4 and Remark 2 of [6])
For the JSQ system described in Theorem 1 with $\rho=\frac{\lambda}{2\mu}<1$, we have
$\mathbb{E}_{\pi}[q_{1}+q_{2}]\leq\frac{\rho(2-\rho)}{1-\rho}.$
Proof of Lemma 1.
We consider a JSQ system with queue length vector $\bm{s}$, driven by the same arrival and potential service processes as $\bm{p}$ and $\bm{q}$ with $s_{j}(0)=\max\left\{q_{j}(0),p_{j}(0)\right\},j\in\{1,2\}$. We couple the three systems $\bm{p},\bm{q},\bm{s}$ similarly to Section 3.1. Formally, let $\bm{s}$ evolve according to
$\begin{split}ds_{1}(t)&=\mathbb{I}_{\left\{s_{1}(t-)\leq s_{2}(t-)\right\}}\mathcal{N}_{a}(dt)-\mathbb{I}_{\left\{s_{1}(t-)\geq 1\right\}}\mathcal{N}_{d}^{1}(dt),\\ ds_{2}(t)&=\mathbb{I}_{\left\{s_{1}(t-)>s_{2}(t-)\right\}}\mathcal{N}_{a}(dt)-\mathbb{I}_{\left\{s_{2}(t-)\geq 1\right\}}\mathcal{N}_{d}^{2}(dt).\end{split}$
Since $\bm{s}(0)\geq\bm{p}(0)$ and $\bm{s}(0)\geq\bm{q}(0)$ componentwise, Lemma 2 (applied to the coupled pairs $(\bm{s},\bm{p})$ and $(\bm{s},\bm{q})$) implies that for all $t\geq 0$, $\bm{s}(t)\geq\bm{p}(t)$ and $\bm{s}(t)\geq\bm{q}(t)$ componentwise. Hence, for any $t\geq 0$, $\bm{s}(t)=(0,0)$ implies $\bm{p}(t)=(0,0)$ and $\bm{q}(t)=(0,0)$. From Lemma 3, the time $\bm{s}$ hits 0, denoted by $T^{\prime}$, is stochastically dominated by $\sum_{i=1}^{s_{1}(0)+s_{2}(0)}T_{0i}$, where $T_{0i}$ are i.i.d. copies of $T_{0}$ (3.1). Therefore,
$\mathbb{E}[T]$ $\leq\mathbb{E}[T^{\prime}]=\mathbb{E}_{\bm{s}(0)}[\mathbb{E}[T^{\prime}]\mid\bm{s}(0)]$
$\leq\mathbb{E}_{\bm{s}(0)}[(s_{1}(0)+s_{2}(0))\mathbb{E}[T_{0}]]=\mathbb{E}_{\bm{q}(0)}\Big[\max\left\{q_{1}(0),p_{1}(0)\right\}+\max\left\{q_{2}(0),p_{2}(0)\right\}\Big]\mathbb{E}[T_{0}].$ (3.5)
Using $\max\left\{x,y\right\}\leq x+y$ for $x,y\geq 0$, (3.5) implies
$\mathbb{E}[T]\leq\mathbb{E}[T_{0}]\Big((x_{1}+x_{2})+\mathbb{E}_{\pi}[q_{1}+q_{2}]\Big).$ (3.6)
Using $\mathbb{E}_{\pi}[q_{1}+q_{2}]\leq\frac{\rho(2-\rho)}{1-\rho}$ from Proposition 2 and recalling $\mathbb{E}[T_{0}]\leq\frac{1}{\lambda}\frac{\rho(2-\rho)}{(1-\rho)^{2}}=\frac{2-\rho}{2\mu(1-\rho)^{2}}$ from (3.4) in Lemma 4, we arrive at $\mathbb{E}[T]\leq C_{1}(\lambda,\mu,\bm{x})$ (2.2). Applying Markov’s inequality $\mathbb{P}\left(T\geq t\right)\leq\frac{\mathbb{E}[T]}{t}$ for $t>0$, we obtain $\mathbb{P}\left(T\geq t\right)\leq\frac{1}{t}C_{1}(\lambda,\mu,\bm{x})$ and we complete the proof of Lemma 1. ∎
3.4 Proof of Lemma 2
Consider the processes $\mathcal{N}_{a}(t),\mathcal{N}_{d}^{1}(t)$, and $\mathcal{N}_{d}^{2}(t)$ in (3.2). Fix a sample path. Let $t_{1}$ be the first time that either an arrival or a potential service completion event occurs. In words, $t_{1}$ is the smallest time $t$ such that $\mathcal{N}_{a}(t)=1$ or $\mathcal{N}_{d}^{1}(t)=1$ or $\mathcal{N}_{d}^{2}(t)=1$. We will show that $p_{i}(t_{1})\geq q_{i}(t_{1})$.
  1. 1.
    If the first event is a potential service completion, then as we assume the two systems are coupled, we have $dp_{j}(t_{1})=-\mathbb{I}_{\left\{p_{j}(0)\geq 1\right\}},dq_{j}(t_{1})=-\mathbb{I}_{\left\{q_{j}(0)\geq 1\right\}}$, where $j$ is the index such that $\mathcal{N}_{d}^{j}(t_{1})=1$. It is easy to verify that $p_{i}(t_{1})\geq q_{i}(t_{1})$ for $i\in\{1,2\}$ given $p_{i}(0)\geq q_{i}(0)$.
  2. 2.
    Next, we assume that the first event is an arrival. For the sake of contradiction, suppose for some $j\in\{1,2\}$, we have $p_{j}(t_{1})<q_{j}(t_{1})$.
    1. (a)
      First, we assume $j=1$. Since at each time there is at most one arrival with probability one, $p_{j}(0)\geq q_{j}(0)$ and $p_{j}(t_{1})<q_{j}(t_{1})$ imply
      $p_{1}(0)=q_{1}(0)$ (3.7)
      and
      $dp_{2}(t_{1})=1,dq_{1}(t_{1})=1.$
      Combining the above equation and the definition of the JSQ policy, we arrive at $p_{1}(0)>p_{2}(0)$ and $q_{2}(0)\geq q_{1}(0)$. Combining this and the assumption $p_{2}(0)\geq q_{2}(0)$, we have
      $p_{1}(0)>p_{2}(0)\geq q_{2}(0)\geq q_{1}(0),$
      which contradicts (3.7).
    2. (b)
      Next, we assume that $j=2$, so $p_{j}(t_{1})<q_{j}(t_{1})$ implies
      $p_{2}(0)=q_{2}(0)$ (3.8)
      and
      $dp_{1}(t_{1})=1,dq_{2}(t_{1})=1.$
      Combining the above equation, $p_{1}(0)\geq q_{1}(0)$, and the definition of the JSQ policy, we arrive at
      $p_{2}(0)\geq p_{1}(0)\geq q_{1}(0)>q_{2}(0),$
      which contradicts (3.8).
Combining the above two cases, we show $p_{i}(t_{1})\geq q_{i}(t_{1})$ for $i\in\{1,2\}$. Proceeding in the same way, we show that $p_{i}(t)\geq q_{i}(t)$ for any $t\geq 0$ and $i\in\{1,2\}$.
Next, we will show that $r(t)\leq r(0)$ holds for every $t\geq 0$. We fix a sample path. Consider the processes $\mathcal{N}_{a}(t),\mathcal{N}_{d}^{1}(t),$ and $\mathcal{N}_{d}^{2}(t)$ in (3.2). Let $t$ denote the first time that either an arrival or a potential service completion event occurs.
  1. 1.
    If the first event is an arrival, then it is obvious that $r(t)=r(0)$.
  2. 2.
    If the first event is a potential service completion, then we let $i$ denote the index such that $\mathcal{N}_{d}^{i}(t)=1$. Since $\bm{p}(0)\geq\bm{q}(0)$, we have $p_{i}(t)\geq q_{i}(t)$ as shown above. Thus, there are two cases:
    1. (a)
      $\mathbb{I}_{\left\{p_{i}(t-)>0\right\}}=\mathbb{I}_{\left\{q_{i}(t-)>0\right\}}$, which implies $r(t)=r(0)$.
    2. (b)
      $p_{i}(t-)>0,q_{i}(t-)=0$, and $dp_{i}(t)=-1,dq_{i}(t)=0$, which implies that $r(t)\leq r(0)$.
Combining the two cases, we show $r(t)\leq r(0)$ for the $t$ defined above. Using the same argument, we show $r(t)\leq r(0)$ holds for every $t\geq 0$.
3.5 Proof of Lemma 3
Consider two coupled JSQ systems with queue length processes $\bm{p}(t)$ and $\bm{q}(t)$ as governed by (3.2). Suppose $\bm{p}(0)=(x+1,y)$ and $\bm{q}(0)=(x,y)$ for some $x\geq 0,y\geq 0$. Then from Lemma 2, for any $t\geq$0, $\bm{q}(t)=(0,0)$ implies $\bm{p}(t)\in\left\{(1,0),(0,1),(0,0)\right\}$.
Next, we show $T(0,1){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T(1,0){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T_{0}$. We can couple two copies of queues $\bm{q}^{(1)}$ and $\bm{q}^{(2)}$ started from $(1,0)$ and $(0,1)$ by using the same arrival process and the same potential service process for the unique initial job (so that if the job completes first, it completes in queue 1 for $\bm{q}^{(1)}$ and in queue 2 for $\bm{q}^{(2)}$ at the same time). Under this coupling, at the first event time, the two copies will be in the same post-jump state, either both at $(0,0)$ or both at $(1,1)$. Thus, $T(0,1){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T(1,0){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T_{0}$.
Combining $\bm{q}(t)=(0,0)\Rightarrow\bm{p}(t)\in\left\{(1,0),(0,1),(0,0)\right\},\forall t\geq 0$, $\quad T(0,1){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T(1,0){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T_{0}$, and the strong Markov property, we have $\mathcal{T}(x+1,y)\preceq\mathcal{T}(x,y)+T_{0}^{\prime}$, where $T_{0}^{\prime}$ is an independent copy of $T_{0}$. Similarly, we can show that $\mathcal{T}(x,y+1)\preceq\mathcal{T}(x,y)+T_{0}^{\prime}$. Iterating these one-step bounds yields $T(x_{1},x_{2})\preceq\sum_{i=1}^{x_{1}+x_{2}}T_{0i}$.
3.6 Proof of Lemma 4
As discussed, we analyze the expected return time of a CTMC constructed by two $M/M/1$ queues. First, we present a general result for the expected hitting time in terms of the $Q$-matrix of a CTMC.
Proposition 3 (Theorem 3.5.3 of [13])
Let $Q$ be an irreducible $Q$-matrix of a CTMC. The following are equivalent:
(i) the CTMC is positive recurrent;
(ii) $Q$ is non-explosive and has a steady state $\pi$.
Moreover, when (ii) holds, we have $m_{i}=\frac{1}{\pi_{i}\nu_{i}}$ for all $i$, where $\nu_{i}$ is the absolute value of $i$-th diagonal entry of $Q$ and $m_{i}$ is the expected time to return to state $i$ when starting from state $i$ including the holding time in state $i$.
Then we use the following result comparing JSQ and a relevant system with two independent $M/M/1$ queues.
Proposition 4 (Theorem 4 of [18])
Let $(q^{\text{J}}_{1}(t),q^{\text{J}}_{2}(t))$ denote the queue length vector for the JSQ system defined in Theorem 1 and $(q^{\text{M}}_{1}(t),q^{\text{M}}_{2}(t))$ denote the queue length vector for the two independent $M/M/1$ queues with $\bm{q}^{\text{J}}(0)=\bm{q}^{\text{M}}(0)$. Further, we assume both $M/M/1$ queues have an arrival rate of $\frac{\lambda}{2}$ and a service rate of $\mu$. Then, there exists a coupling such that
$q^{\text{J}}_{i}(t)\leq q^{\text{M}}_{i}(t),i\in\{1,2\},t\geq 0\text{ almost surely}.$
This shows that, in this coupling, almost surely, $\bm{q}^{\text{M}}(t)=(0,0)$ implies $\bm{q}^{\text{J}}(t)=(0,0)$ for any $t\geq 0$.
With the above tool, we prove Lemma 4 as follows. Consider two independent $M/M/1$ queues with arrival rates $\lambda/2$ and service rates $\mu$, and denote their queue lengths as $q^{\text{M}}_{1}(t)$ and $q^{\text{M}}_{2}(t)$. Then their Cartesian product, $\bm{q}^{\text{M}}(t)=(q^{\text{M}}_{1}(t),q^{\text{M}}_{2}(t))$, is also a CTMC. Since each marginal queue is irreducible and positive recurrent under the condition $\lambda/2<\mu$, it follows that the joint process $\bm{q}^{\text{M}}(t)$ is also irreducible and positive recurrent, with the stationary distribution given by the product of the two marginal stationary distributions. Under stationarity, $\mathbb{P}\left(q^{\text{M}}_{1}=k\right)=\mathbb{P}\left(q^{\text{M}}_{2}=k\right)=\rho^{k}(1-\rho),\forall k\in\mathbb{Z}_{+}$. Since they are independent, $\mathbb{P}_{\pi^{\text{M}}}\left(q^{\text{M}}_{1}=q^{\text{M}}_{2}=0\right)=(1-\rho)^{2}$, where $\pi^{\text{M}}$ is the stationary distribution of $\bm{q}^{\text{M}}$.
Let $m^{\text{M}}_{(0,0)}$ denote the expected time for $\bm{q}^{\text{M}}$ to return to $(0,0)$ when starting from $(0,0)$ including the holding time in $(0,0)$. Applying Proposition 3 to $\bm{q}^{\text{M}}$, and noting that the total transition rate out of state $(0,0)$ is $\nu^{\text{M}}_{(0,0)}=\lambda$, we obtain
$m^{\text{M}}_{(0,0)}=\frac{1}{\nu^{\text{M}}_{(0,0)}\mathbb{P}_{\pi^{\text{M}}}\left(q^{\text{M}}_{1}=q^{\text{M}}_{2}=0\right)}=\frac{1}{\lambda(1-\rho)^{2}}.$ (3.9)
Starting from state $(0,0)$, the process $\bm{q}^{\text{M}}$ leaves $(0,0)$ after an $\mathsf{Exp}(\lambda)$ time and moves to either $(1,0)$ or $(0,1)$. Let $T_{(1,0)}^{(0,0)}$ be the time for $\bm{q}^{\text{M}}$ to hit state $(0,0)$ starting from state $(1,0)$. By symmetry,
$m^{\text{M}}_{(0,0)}=\frac{1}{\lambda}+\mathbb{E}\left[T^{(0,0)}_{(1,0)}\right].$
Combining the above equation with (3.9), we have
$\mathbb{E}\left[T^{(0,0)}_{(1,0)}\right]=\frac{1}{\lambda}\left(\frac{1}{(1-\rho)^{2}}-1\right)=\frac{1}{\lambda}\frac{\rho(2-\rho)}{(1-\rho)^{2}}.$ (3.10)
Next, consider a two-queue JSQ system $\bm{q}^{\text{J}}(t)=(q_{1}(t),q_{2}(t))$ (with an arrival rate $\lambda$ and each queue has a service rate $\mu$) and the product chain $\bm{q}^{\text{M}}(t)=(q^{\text{M}}_{1}(t),q^{\text{M}}_{2}(t))$ of two independent $M/M/1$ queues defined previously, started from the same initial state $\bm{q}^{\text{J}}(0)=\bm{q}^{\text{M}}(0)=(1,0)$. By Proposition 4, there exists a copuling such that
$T_{0}=\inf\left\{t\geq 0:\bm{q}^{\text{J}}(t)=(0,0)\right\}\leq\inf\left\{t\geq 0:\bm{q}^{\text{M}}(t)=(0,0)\right\}=T^{(0,0)}_{(1,0)}\qquad\text{almost surely.}$
Using the above inequality and (3.10), we show $\mathbb{E}[T_{0}]\leq\mathbb{E}\left[T^{(0,0)}_{(1,0)}\right]=\frac{1}{\lambda}\frac{\rho(2-\rho)}{(1-\rho)^{2}}$ and we finish the proof.
3.7 Complete proof of Theorem 1
To prove Theorem 1, we will construct a tighter upper bound on $\mathbb{E}[T_{0}]$ in Lemma 5. Throughout, when we write $Z,Z^{\prime}$, we mean random variables that are independent copies of $Z$, and are independent of all random variables defined previously (unless stated otherwise).
First, we define $T_{1}$ as
$T_{1}=$ $\inf\left\{t\geq 0:\bm{q}(t)\in\left\{(0,1),(1,0)\right\}\mid\bm{q}(0)=(1,1)\right\}.$ (3.11)
Lemma 5
Define two independent random variables $X\sim\mathsf{Exp}\left(\lambda+2\mu\right)$ and $Y\sim\mathsf{Exp}\left(\lambda+\mu\right)$, then $T_{1}\preceq A$ and $T_{0}{\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}B$ where $A$ and $B$ are random variables defined as
[…]
and
[…]
where $T_{1}^{\prime}{\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T_{1},T_{0}^{\prime}{\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T_{0}$, $T_{1}^{\prime},T_{0}^{\prime}$ are independent, and independent of $T_{1},T_{0},X,Y$. Furthermore, if $\rho<\frac{1}{\sqrt{2}}$,
$\left\{\begin{aligned} &\mathbb{E}[T_{1}]\leq&\frac{\lambda+\mu}{2\mu^{2}-\lambda^{2}},\\ &\mathbb{E}[T_{0}]\leq&\frac{\lambda+2\mu}{2\mu^{2}-\lambda^{2}}.\end{aligned}\right$. (3.14)
Further,
$\mathbb{E}[T_{0}]\leq\ \frac{\lambda+2\mu}{2\mu^{2}-\lambda^{2}}=\frac{1}{\lambda}\frac{1+\frac{2\mu}{\lambda}}{2\frac{\mu^{2}}{\lambda^{2}}-1}=\frac{1}{\lambda}\frac{1+\frac{1}{\rho}}{\frac{1}{2}\frac{1}{\rho^{2}}-1}=\frac{1}{\lambda}\frac{2\rho^{2}+2\rho}{1-2\rho^{2}}.$ (3.15)
We replace (3.4) in the proof of Lemma 1 by the sharper bound (3.15) from Lemma 5 (and taking the minimum with (3.4) when appropriate). Recalling the expression of $K(\rho)$ in (2.3), since
$\frac{\rho(2-\rho)}{(1-\rho)^{2}}\cdot\left(\frac{2+2\rho}{1-2\rho^{2}}\cdot\frac{(1-\rho)^{2}}{2-\rho}\right)=\frac{2\rho^{2}+2\rho}{1-2\rho^{2}},$
we show $\mathbb{P}\left(T\geq t\right)\leq\frac{1}{t}C^{K}_{1}(\lambda,\mu,\bm{x}),$ where $C^{K}_{1}(\lambda,\mu,\bm{x})$ is defined in (2.4). Thus, we complete the proof of Theorem 1.
3.8 Proof of Lemma 5
To prove Lemma 5, we use Lemma 2 to construct a recursive relationship between $T_{0}$ (3.1) and $T_{1}$ (3.11). We define
$\mathcal{T}(\bm{x};\mathcal{A}):=\inf\left\{t\geq 0:\bm{p}(t)\in\mathcal{A}\mid\bm{p}(0)=\bm{x}\right\}$
as the random time it takes for a JSQ queue length process $\bm{p}(t)$ to hit a set $\mathcal{A}$ starting from initial queue length $\bm{x}$. We analyze $T_{1}$ (3.11) and $T_{0}$ (3.1) separately.
  1. 1.
    We first analyze $T_{1}$ (3.11). We let $\bm{p}(0)=(1,1)$, and let $t_{1}$ denote the first time that there will be either an arrival or a service completion event. Then $t_{1}{\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}X$.
    1. (a)
      If the first event is a service completion, which has probability $\frac{2\mu}{\lambda+2\mu}$, then $\bm{p}(t_{1})\in\left\{(1,0),(0,1)\right\}$.
    2. (b)
      If the first event is an arrival, which has probability $\frac{\lambda}{\lambda+2\mu}$, then $\bm{p}(t_{1})=(2,1)$. Now consider two JSQ systems with queue length vectors $\bm{s}(t),\bm{w}(t)$ and $\bm{s}(0)=(2,1),\bm{w}(0)=(1,1)$. We also assume they are coupled with arrival and service, as defined in Section 3.1. Thus, $\bm{w}$ will hit the set $\left\{(1,0),(0,1)\right\}$ at some time $t_{2}{\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T_{1}$ with $T_{1}$ defined in (3.11). By Lemma 2,
      $\bm{s}(t_{2})\in\left\{(1,1),(1,0),(0,1),(2,0),(0,2)\right\}.$
      Next, we analyze the total time for $\bm{s}$ to hit $\left\{(1,0),(0,1)\right\}$ starting from its initial state $\bm{s}(0)=(2,1)$, which can be written as $\mathcal{T}((2,1);\left\{(1,0),(0,1)\right\})$. There are several cases:
      1. i.
        If $\bm{s}(t_{2})\in\left\{(1,0),(0,1)\right\}$, then we have $\mathcal{T}((2,1);\left\{(1,0),(0,1)\right\})=t_{2}$ conditioned on this event.
      2. ii.
        If $\bm{s}(t_{2})=(1,1)$, and suppose it takes another $T_{1}^{\prime}$ time for $\bm{s}$ to hit $\left\{(1,0),(0,1)\right\}$. In addition, using Lemma 2, we can show $T_{1}^{\prime}\preceq T_{0}^{\prime}$: indeed, couple $\bm{s}$ (started from $(1,1)$) with a JSQ process $\bm{q}$ started from $(1,0)$ using the common arrival/service construction, and let
        $\tau:=\inf\{t\geq 0:\bm{q}(t)=(0,0)\}\stackrel{{\scriptstyle d}}{{=}}T_{0}^{\prime}.$
        By Lemma 2, $\bm{q}(\tau)=(0,0)$ implies $\bm{s}(\tau)\in\left\{(0,0),(1,0),(0,1)\right\}$, so $\bm{s}$ must have hit $\left\{(1,0),(0,1)\right\}$ by time $\tau$. Thus, $T_{1}^{\prime}\leq\tau$ almost surely, which yields $T_{1}^{\prime}\preceq T_{0}^{\prime}$.
      3. iii.
        If $\bm{s}(t_{2})\in\left\{(2,0),(0,2)\right\}$. Couple a JSQ process $\bm{p}$ with $\bm{p}(0)=(2,0)$ and another JSQ process $\bm{q}$ with $\bm{q}(0)=(1,0)$ using the common arrival/service construction, and let $\tau:=\inf\{t\geq 0:\bm{q}(t)=(0,0)\}\overset{d}{=}T_{0}$. By Lemma 2, $\bm{q}(\tau)=(0,0)$ implies $\bm{p}(\tau)\in\{(0,0),(1,0),(0,1)\}$; moreover, any sample path from $(2,0)$ to $(0,0)$ must hit $(1,0)$ or $(0,1)$ beforehand. Thus, $\mathcal{T}((2,0);\left\{(1,0),(0,1)\right\})\leq\tau$ a.s., and therefore $\mathcal{T}((2,0);\left\{(1,0),(0,1)\right\})\preceq T_{0}$ (the case $s(t_{2})=(0,2)$ follows by symmetry).
      In summary, the above argument shows that
      $\mathcal{T}((2,1);\left\{(1,0),(0,1)\right\})\preceq T_{1}+T_{0}.$
      Therefore, $T_{1}\preceq A$ with $A$ defined in (3.12).
  2. 2.
    Now we analyze $T_{0}$ (3.1). Let $\bm{p}(0)=(1,0)$, and let $t_{1}$ denote the first time at which either an arrival or a service completion event occurs. Then $t_{1}{\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}Y$.
    • If the first event is a service completion, which has probability $\frac{\mu}{\lambda+\mu}$, then $\bm{p}(t_{1})=(0,0)$.
    • If the first event is an arrival, which has probability $\frac{\lambda}{\lambda+\mu}$, then $\bm{p}(t_{1})=(1,1)$, and the time it takes for the queue length vector $\bm{p}(t)$ to hit $(0,0)$, which we denote as $\mathcal{T}((1,1);(0,0))$, satisfies
      $\mathcal{T}((1,1);(0,0)){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T_{1}^{\prime}+T_{0}^{\prime}.$
      Thus, we have shown $T_{0}{\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}B$ with $B$ defined in (3.13).
Combining $T_{1}\preceq A$, and $T_{0}{\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}B$, we arrive at
$\left\{\begin{aligned} &\mathbb{E}[T_{1}]\leq&\frac{1}{\lambda+2\mu}+\frac{\lambda}{\lambda+2\mu}\left(\mathbb{E}[T_{1}]+\mathbb{E}[T_{0}]\right),\\ &\mathbb{E}[T_{0}]=&\frac{1}{\lambda+\mu}+\frac{\lambda}{\lambda+\mu}\left(\mathbb{E}[T_{1}]+\mathbb{E}[T_{0}]\right).\end{aligned}\right$.
It follows that $2\mu\mathbb{E}[T_{1}]\leq 1+\lambda\mathbb{E}[T_{0}]=1+\frac{\lambda}{\mu}+\lambda\frac{\lambda}{\mu}\mathbb{E}[T_{1}]$ and we know this system is positive recurrent, which implies that $\mathbb{E}[T_{1}]<\infty$ and $\mathbb{E}[T_{0}]<\infty$. Solving the above inequality yields (3.14) if $\rho<\frac{1}{\sqrt{2}}$.
3.9 Proof of Corollary 1
To show Corollary 1, we will derive an upper bound for $\mathbb{E}[q_{1}(t)^{2}+q_{2}(t)^{2}]$, where $\bm{q}(t)$ is the queue length process for a JSQ system. To do this, due to Proposition 4, it suffices to analyze two independent $M/M/1$ systems. Thus, in Lemma 6, we analyze the corresponding $M/M/1$ systems.
Lemma 6
Consider an $M/M/1$ queue with an arrival rate $\lambda$ and a service rate $\mu$. Let $q(t)$ be its queue length at time $t$ and assume $\rho=\frac{\lambda}{\mu}<1$. Let $\pi$ denote its steady state. Then
$\mathbb{E}_{\pi}[q^{2}]=\frac{\rho(1+\rho)}{(1-\rho)^{2}},$
and
$\mathbb{E}[q(t)^{2}|q(0)=x]\leq x^{2}+2x\frac{\rho}{1-\rho}+\frac{\rho(1+\rho)}{(1-\rho)^{2}}.$
Proof of Lemma 6.
The first equality follows from the fact that the steady-state distribution of an $M/M/1$ system is $\pi(k)=\rho^{k}(1-\rho)$ so $\mathbb{E}_{\pi}[q^{2}]=\sum_{k=0}^{\infty}k^{2}\rho^{k}(1-\rho)=\frac{\rho(1+\rho)}{(1-\rho)^{2}}$.
Let $q_{x}(t)$ denote the queue length process starting from $q(0)=x\geq 0$, and let $q_{0}(t)$ denote the queue length process starting from $q(0)=0$. We couple the two processes by using the same arrival and potential service processes: both queues receive arrivals from a common Poisson process of rate $\lambda$ and services from a common Poisson process of rate $\mu$. Under this coupling, the process starting from $x$ remains below the one starting from 0 plus $x$, that is,
$q_{x}(t)\leq q_{0}(t)+x\text{ almost surely}\text{ for all }t\geq 0.$
Squaring both sides and taking expectations, we obtain
$\mathbb{E}[q_{x}(t)^{2}]$ $\leq\mathbb{E}[(q_{0}(t)+x)^{2}]$
$=\mathbb{E}[q_{0}(t)^{2}]+2x\,\mathbb{E}[q_{0}(t)]+x^{2}.$
Thus,
$\mathbb{E}[q(t)^{2}\mid q(0)=x]$ $=\mathbb{E}[q_{x}(t)^{2}]$
$\leq\mathbb{E}[q_{0}(t)^{2}]+2x\,\mathbb{E}[q_{0}(t)]+x^{2}$
$=\mathbb{E}[q(t)^{2}\mid q(0)=0]+2x\,\mathbb{E}[q(t)\mid q(0)=0]+x^{2}.$ (3.16)
In addition, let $q_{\pi}(t)$ denote the queue length process starting from $q(0)\sim\pi$. We couple the two processes by using the same arrival and service processes. Since $q_{0}(0)\leq q_{\pi}(0)$ almost surely,
$q_{0}(t)\leq q_{\pi}(t)\text{ almost surely}\text{ for all }t\geq 0.$
Thus,
$\mathbb{E}[q(t)^{2}\mid q(0)=0]$ $\leq\mathbb{E}_{\pi}[q^{2}]=\frac{\rho(1+\rho)}{(1-\rho)^{2}},$ (3.17)
$\mathbb{E}[q(t)\mid q(0)=0]$ $\leq\mathbb{E}_{\pi}[q]=\frac{\rho}{1-\rho}.$ (3.18)
Combining (3.17), (3.18), and (3.16), we have
$\mathbb{E}[q(t)^{2}|q(0)=x]\leq x^{2}+2x\frac{\rho}{1-\rho}+\frac{\rho(1+\rho)}{(1-\rho)^{2}},$
and we finish the proof. ∎
Now we are ready to prove Corollary 1.
Proof of Corollary 1.
Fix an integer $R>0$. Then
$\begin{split}&|\mathbb{E}[q_{1}(t)]-\mathbb{E}_{\pi}[q_{1}]|\\ &=|\mathbb{E}[q_{1}(t)\mathbb{I}_{\left\{q_{1}(t)\leq R\right\}}]+\mathbb{E}[q_{1}(t)\mathbb{I}_{\left\{q_{1}(t)>R\right\}}]-\mathbb{E}_{\pi}[q_{1}\mathbb{I}_{\left\{q_{1}\leq R\right\}}]-\mathbb{E}_{\pi}[q_{1}\mathbb{I}_{\left\{q_{1}>R\right\}}]|\\ &\leq|\mathbb{E}[q_{1}(t)\mathbb{I}_{\left\{q_{1}(t)\leq R\right\}}]-\mathbb{E}_{\pi}[q_{1}\mathbb{I}_{\left\{q_{1}\leq R\right\}}]|+|\mathbb{E}[q_{1}(t)\mathbb{I}_{\left\{q_{1}(t)>R\right\}}]-\mathbb{E}_{\pi}[q_{1}\mathbb{I}_{\left\{q_{1}>R\right\}}]|\\ &\stackrel{{\scriptstyle(a)}}{{=}}\left|\sum_{m=0}^{R-1}\mathbb{P}\left(m<q_{1}(t)\leq R\right)-\mathbb{P}_{\pi}\left(m<q_{1}\leq R\right)\right|+|\mathbb{E}[q_{1}(t)\mathbb{I}_{\left\{q_{1}(t)>R\right\}}]-\mathbb{E}_{\pi}[q_{1}\mathbb{I}_{\left\{q_{1}>R\right\}}]|\\ &\stackrel{{\scriptstyle(b)}}{{\leq}}\sum_{m=0}^{R-1}|\mathbb{P}\left(m<q_{1}(t)\leq R\right)-\mathbb{P}_{\pi}\left(m<q_{1}\leq R\right)|+\frac{1}{R}\Big(\mathbb{E}[q_{1}(t)^{2}\mathbb{I}_{\left\{q_{1}(t)>R\right\}}]+\mathbb{E}_{\pi}[(q_{1})^{2}\mathbb{I}_{\left\{q_{1}>R\right\}}]\Big)\\ &\stackrel{{\scriptstyle(c)}}{{\leq}}R\cdot d_{\mathrm{TV}}\left(\mathbb{P}_{\bm{x}}^{t},\pi\right)+\frac{1}{R}\Big(\mathbb{E}[q_{1}(t)^{2}]+\mathbb{E}_{\pi}[(q_{1})^{2}]\Big).\end{split}$ (3.19)
In (3.19), $(a)$ follows from the equality $\mathbb{E}[X\mathbb{I}_{\left\{X\leq a\right\}}]=\sum_{m=0}^{a-1}\mathbb{P}\left(m<X\leq a\right)$ for any non-negative integer-valued random variable $X$ and $a\in\mathbb{N}_{+}$, $(b)$ follows from the triangle inequality, the fact that $x\mathbb{I}_{\left\{x\geq R\right\}}\leq\frac{x^{2}}{R}\mathbb{I}_{\left\{x\geq R\right\}}$ for any $x\geq 0$, and $|x-y|\leq x+y$ for any $x,y\geq 0$, and $(c)$ follows from $|\mathbb{P}(A)-\mathbb{Q}(A)|\leq d_{\mathrm{TV}}\left(\mathbb{P},\mathbb{Q}\right)$ with $A=\left\{(q_{1},q_{2}):m<q_{1}\leq R\right\}$. Similar to (3.19), we have
$|\mathbb{E}[q_{2}(t)]-\mathbb{E}_{\pi}[q_{2}]|\leq Rd_{\mathrm{TV}}\left(\mathbb{P}_{\bm{x}}^{t},\pi\right)+\frac{1}{R}\Big(\mathbb{E}[q_{2}(t)^{2}]+\mathbb{E}_{\pi}[(q_{2})^{2}]\Big).$
The triangle inequality yields
$|\mathbb{E}[q_{1}(t)+q_{2}(t)]-\mathbb{E}_{\pi}[q_{1}+q_{2}]|\leq 2Rd_{\mathrm{TV}}\left(\mathbb{P}_{\bm{x}}^{t},\pi\right)+\frac{1}{R}\Big(\mathbb{E}[q_{1}(t)^{2}+q_{2}(t)^{2}]+\mathbb{E}_{\pi}[(q_{1})^{2}+(q_{2})^{2}]\Big).$
Since $\rho=\frac{\lambda}{2\mu}$, then by Lemma 6 and Proposition 4, $\mathbb{E}_{\pi}[(q_{1})^{2}+(q_{2})^{2}]\leq\frac{2\rho(1+\rho)}{(1-\rho)^{2}}$, and
$\mathbb{E}[q_{1}(t)^{2}+q_{2}(t)^{2}]\leq(x_{1}^{2}+x_{2}^{2})+2(x_{1}+x_{2})\frac{\rho}{1-\rho}+\frac{2\rho(1+\rho)}{(1-\rho)^{2}}.$
Taking $R=\lceil\sqrt{t}\rceil\leq\sqrt{t}+1$ yields
$|\mathbb{E}[q_{1}(t)+q_{2}(t)]-\mathbb{E}_{\pi}[q_{1}+q_{2}]|\leq 2(\sqrt{t}+1)d_{\mathrm{TV}}\left(\mathbb{P}_{\bm{x}}^{t},\pi\right)+\frac{1}{\sqrt{t}}C_{2}(\lambda,\mu,\bm{x}).$
By Theorem 1, $d_{\mathrm{TV}}\left(\mathbb{P}_{\bm{x}}^{t},\pi\right)\leq\frac{1}{t}C_{1}^{K}(\lambda,\mu,\bm{x})$, where $C_{1}^{K}$ is defined in (2.4).
Therefore, we conclude that
$|\mathbb{E}[q_{1}(t)+q_{2}(t)]-\mathbb{E}_{\pi}[q_{1}+q_{2}]|\leq\frac{1}{\sqrt{t}}\Big(2C_{1}^{K}(\lambda,\mu,\bm{x})+C_{2}(\lambda,\mu,\bm{x})\Big)+\frac{2}{t}C_{1}^{K}(\lambda,\mu,\bm{x}).$
Using Proposition 2, we finish proving (2.7). ∎
4 Conclusion and future work
In this paper, we present a convergence rate result for a load balancing system with two queues under the JSQ policy, which is valid for any $\rho\in(0,1)$. Our method is based on a novel and simple coupling construction and on computing an upper bound on the expected hitting time. Instead of relying on approximations of the stochastic systems in relevant limits, we directly analyze the original system, yielding clean insights regarding the convergence rate and how it depends on $\rho$.
Below, we discuss some interesting future directions. Recall that we have established a convergence rate that is of order $O\left(\frac{1}{(1-\rho)^{3}}\frac{1}{t}\right)$. However, based on the results for a single-server queue [14] and also based on the scaling used in diffusion approximations, one expects an exponential convergence rate of the form $O\left(e^{-\alpha\epsilon^{2}t}\right)$ with $\epsilon=1-\rho$ for an appropriate constant $\alpha$. Establishing such a strong bound on convergence is a future research direction, and we believe it requires adopting novel methodological approaches, particularly by exploiting state space collapse results from the literature. Another possible approach is to compute an upper bound on the moment generating function of the hitting time. An alternative direction is to extend our analysis to a system with $n\geq 3$ servers and heterogeneous service rates. We also hope the technique we have developed in this paper can further spur analysis of more complicated stochastic systems.
Acknowledgments
This work was partially supported by NSF grants EPCN-2144316 and CMMI-2140534.

### v4

@@
 Efficient load balancing is central to optimizing resource allocation and minimizing job delays in modern service systems and data centers.
-The Join–the-Shortest-Queue (JSQ) policy is a popular load balancing algorithm due to its simplicity and optimality under various criteria.
+The Join–the-Shortest-Queue (JSQ) policy is a popular load balancing algorithm due to its simplicity and strong performance under various criteria.
 Since its introduction by Winston [20], it has received significant attention over the decades.
@@
 Further, researchers typically conduct analysis in some asymptotic regimes, such as the heavy-traffic regime.
-For example, there are various bounds for the queue length of a JSQ system in its steady state, such as [15, Theorem 10.2.3].
+For example, there are various bounds on the queue length of a JSQ system in its steady state, such as [15, Theorem 10.2.3].
 However, traffic patterns in modern data centers tend to be time-varying [2].
-This necessitates careful finite-time analysis of JSQ performance, rather than relying only on steady-state behavior.
+This necessitates a careful finite-time analysis of JSQ performance, rather than relying only on steady-state behavior.
 Motivated by this, we focus on analyzing the transient behavior of a JSQ system.
@@
 However, unsurprisingly, simulation results in Figure 1 reveal that the convergence rate of a JSQ system can be quite slow when the traffic intensity $\rho$ approaches 1.
-Thus, to conduct more careful analysis of the JSQ systems, we need to understand their transient behavior.
-In addition, prior results are not applicable to light-traffic settings since they rely on state space collapse (SSC), which is only associated with the heavy-traffic regime.
+Thus, to conduct more careful analysis of the JSQ systems, we need to understand their transient behavior.Many prior results rely on state space collapse phenomena in heavy traffic and do not directly yield explicit finite-time bounds for light traffic.
 In this paper, we address this challenge and obtain an explicit upper bound on the difference between the transient distribution of a JSQ system and its steady state for any traffic intensity $\rho\in(0,1)$.
@@
 The pair of plots clearly shows that the convergence rates of JSQ systems heavily depend on $\rho$.
-We focus on analyzing the distribution $\mathbb{P}^{t}_{\bm{x}}$ of the queue length vector at time $t$, for a JSQ system with initial queue length $\bm{x}$.
+We focus on analyzing the distribution $\mathbb{P}^{t}_{\bm{x}}$ of the queue length vector at time $t$, for a JSQ system with an initial queue length vector $\bm{x}$.
 In particular, we consider its total variation (TV) distance to the steady state $\pi$ at every time $t\geq 0$.
@@
 For general continuous-time Markov chains (CTMCs), analyzing the hitting time usually requires extensive exploitation of the structure of the CTMC [4].
-We use several structures of the JSQ system, and show that the problem can be reduced to providing an upper bound on $\mathbb{E}[T_{0}]$, where $T_{0}$, to be defined in (3.1), is the time it takes for a JSQ system’s queue length vector to hit the origin starting from state $(1,0)$.
-After upper bounding $\mathbb{E}[T_{0}]$, we use Markov’s inequality to arrive at a bound on $d_{\mathrm{TV}}\left(\mathbb{P}^{t}_{\bm{x}},\pi\right)$.
+We use several structural properties of the JSQ system, and show that the problem can be reduced to providing an upper bound on $\mathbb{E}[T_{0}]$, where $T_{0}$, to be defined in (3.1), is the time it takes for a JSQ system’s queue length vector to hit the origin starting from state $(1,0)$.
 1.1 Main contribution
@@
 In this work, we analyze the transient behavior through a hitting time analysis (see, e.g., [10]).
-Though there are some results related to hitting times or first passage times for JSQ systems [21, 17], they cannot directly be used for our problem.
+Though there are some results related to hitting times or first passage times for JSQ systems [21, 17], they cannot be used directly for our problem.
 For a simple system such as an $M/M/1$ system, one can obtain the moment generating function (MGF) of the hitting time for the queue length to hit state 0 using a construction of an exponential martingale and the optional stopping theorem [14, Proposition 5.4].
@@
 However, this property is not true for a JSQ system.
-For example, consider a JSQ system with an initial queue length $(x,y)$ with $x,y\geq 1$.
+For example, consider a JSQ system with an initial queue length $(x,y)$, where $x,y\geq 1$.
 It is possible that the queue length vector never hits state (x1,y1)(x-1,y-1) before hitting $(0,0)$.
@@
 Corollary 1
-Consider a JSQ system as described in Theorem 1 with $\rho=\frac{\lambda}{2\mu}<1$ and initial queue length vector $\bm{x}$. Then
+Consider a JSQ system as described in Theorem 1 with $\rho=\frac{\lambda}{2\mu}<1$ and an initial queue length vector $\bm{x}$. Then
 $|\mathbb{E}[q_{1}(t)+q_{2}(t)]-\mathbb{E}_{\pi}[q_{1}+q_{2}]|\leq\frac{1}{\sqrt{t}}\Big(2C_{1}^{K}(\lambda,\mu,\bm{x})+C_{2}(\lambda,\mu,\bm{x})\Big)+\frac{2}{t}C_{1}^{K}(\lambda,\mu,\bm{x})$ (2.6)
@@
 Proof of Lemma 1.
-We consider a JSQ system with queue length vector $\bm{s}$, driven by the same arrival and potential service processes as $\bm{p}$ and $\bm{q}$ with $s_{j}(0)=\max\left\{q_{j}(0),p_{j}(0)\right\},j\in\{1,2\}$.
+We consider a JSQ system with queue length vector $\bm{s}$, driven by the same arrival and potential service processes as $\bm{p}$ and $\bm{q}$, with $s_{j}(0)=\max\left\{q_{j}(0),p_{j}(0)\right\},j\in\{1,2\}$.
 We couple the three systems $\bm{p},\bm{q},\bm{s}$ similarly to Section 3.1.
@@
 $\begin{split}ds_{1}(t)&=\mathbb{I}_{\left\{s_{1}(t-)\leq s_{2}(t-)\right\}}\mathcal{N}_{a}(dt)-\mathbb{I}_{\left\{s_{1}(t-)\geq 1\right\}}\mathcal{N}_{d}^{1}(dt),\\ ds_{2}(t)&=\mathbb{I}_{\left\{s_{1}(t-)>s_{2}(t-)\right\}}\mathcal{N}_{a}(dt)-\mathbb{I}_{\left\{s_{2}(t-)\geq 1\right\}}\mathcal{N}_{d}^{2}(dt).\end{split}$
-Since $\bm{s}(0)\geq\bm{p}(0)$ and $\bm{s}(0)\geq\bm{q}(0)$ componentwise, Lemma 2 (applied to the coupled pairs $(\bm{s},\bm{p})$ and $(\bm{s},\bm{q})$) implies that for all $t\geq 0$, $\bm{s}(t)\geq\bm{p}(t)$ and $\bm{s}(t)\geq\bm{q}(t)$ componentwise.
-Hence, for any $t\geq 0$, $\bm{s}(t)=(0,0)$ implies $\bm{p}(t)=(0,0)$ and $\bm{q}(t)=(0,0)$.
 From Lemma 3, the time $\bm{s}$ hits 0, denoted by $T^{\prime}$, is stochastically dominated by $\sum_{i=1}^{s_{1}(0)+s_{2}(0)}T_{0i}$, where $T_{0i}$ are i.i.d.
 copies of $T_{0}$ (3.1).
-Therefore,
+Since $\bm{s}(0)\geq\bm{p}(0)$ and $\bm{s}(0)\geq\bm{q}(0)$ componentwise, Lemma 2 (applied to the coupled pairs $(\bm{s},\bm{p})$ and $(\bm{s},\bm{q})$) implies that for all $t\geq 0$, $\bm{s}(t)\geq\bm{p}(t)$ and $\bm{s}(t)\geq\bm{q}(t)$ componentwise.
+Thus, for any $t\geq 0$, $\bm{s}(t)=(0,0)$ implies $\bm{p}(t)=(0,0)$ and $\bm{q}(t)=(0,0)$, or $T\leq T^{\prime}$ almost surely.
+Since $T^{\prime}\preceq\sum_{i=1}^{s_{1}(0)+s_{2}(0)}T_{0i}$, we have
 $\mathbb{E}[T]$ $\leq\mathbb{E}[T^{\prime}]=\mathbb{E}_{\bm{s}(0)}[\mathbb{E}[T^{\prime}]\mid\bm{s}(0)]$
 $\leq\mathbb{E}_{\bm{s}(0)}[(s_{1}(0)+s_{2}(0))\mathbb{E}[T_{0}]]=\mathbb{E}_{\bm{q}(0)}\Big[\max\left\{q_{1}(0),p_{1}(0)\right\}+\max\left\{q_{2}(0),p_{2}(0)\right\}\Big]\mathbb{E}[T_{0}].$ (3.5)
-Using $\max\left\{x,y\right\}\leq x+y$ for $x,y\geq 0$, (3.5) implies
+Using $\max\left\{x,y\right\}\leq x+y$ for $x,y\geq 0$ and $\bm{q}(0){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}\pi$, (3.5) implies
 $\mathbb{E}[T]\leq\mathbb{E}[T_{0}]\Big((x_{1}+x_{2})+\mathbb{E}_{\pi}[q_{1}+q_{2}]\Big).$ (3.6)
@@
       and
-      $dp_{2}(t_{1})=1,dq_{1}(t_{1})=1.$
-      Combining the above equation and the definition of the JSQ policy, we arrive at $p_{1}(0)>p_{2}(0)$ and $q_{2}(0)\geq q_{1}(0)$.
-Combining this and the assumption $p_{2}(0)\geq q_{2}(0)$, we have
+      $dp_{2}(t_{1})=1,dq_{1}(t_{1})=1.$ (3.8)
+      Combining (3.8) and the definition of the JSQ policy, we arrive at $p_{1}(0)>p_{2}(0)$ and $q_{2}(0)\geq q_{1}(0)$. Combining this and the assumption $p_{2}(0)\geq q_{2}(0)$, we have
       $p_{1}(0)>p_{2}(0)\geq q_{2}(0)\geq q_{1}(0),$
@@
       and
-      $dp_{1}(t_{1})=1,dq_{2}(t_{1})=1.$
-      Combining the above equation, $p_{1}(0)\geq q_{1}(0)$, and the definition of the JSQ policy, we arrive at
+      $dp_{1}(t_{1})=1,dq_{2}(t_{1})=1.$ (3.10)
+      Combining (3.10), $p_{1}(0)\geq q_{1}(0)$, and the definition of the JSQ policy, we arrive at
       $p_{2}(0)\geq p_{1}(0)\geq q_{1}(0)>q_{2}(0),$
@@
 Suppose $\bm{p}(0)=(x+1,y)$ and $\bm{q}(0)=(x,y)$ for some $x\geq 0,y\geq 0$.
-Then from Lemma 2, for any $t\geq$0, $\bm{q}(t)=(0,0)$ implies $\bm{p}(t)\in\left\{(1,0),(0,1),(0,0)\right\}$.
-Next, we show $T(0,1){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T(1,0){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T_{0}$.
-We can couple two copies of queues $\bm{q}^{(1)}$ and $\bm{q}^{(2)}$ started from $(1,0)$ and $(0,1)$ by using the same arrival process and the same potential service process for the unique initial job (so that if the job completes first, it completes in queue 1 for $\bm{q}^{(1)}$ and in queue 2 for $\bm{q}^{(2)}$ at the same time).
+From Lemma 2, for any $t\geq 0$, $\bm{q}(t)=(0,0)$ implies $\bm{p}(t)\in\left\{(1,0),(0,1),(0,0)\right\}$.
+Next, we show that $T(0,1){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T(1,0){\mathrel{\mathop{\kern 0.0pt=}\limits^{d}}}T_{0}$.
+We can couple two copies of queues $\bm{q}^{(1)}$ and $\bm{q}^{(2)}$ started from $(1,0)$ and $(0,1)$ by using the same arrival process and the same potential service processes for the unique initial job (so that if the job completes first, it completes in queue 1 for $\bm{q}^{(1)}$ and in queue 2 for $\bm{q}^{(2)}$ at the same time).
 Under this coupling, at the first event time, the two copies will be in the same post-jump state, either both at $(0,0)$ or both at $(1,1)$.
@@
 Then, there exists a coupling such that
-$q^{\text{J}}_{i}(t)\leq q^{\text{M}}_{i}(t),i\in\{1,2\},t\geq 0\text{ almost surely}.$
+$\sum_{i=1}^{2}q^{\text{J}}_{i}(t)\leq\sum_{i=1}^{2}q^{\text{M}}_{i}(t),t\geq 0\text{ almost surely}.$
 This shows that, in this coupling, almost surely, $\bm{q}^{\text{M}}(t)=(0,0)$ implies $\bm{q}^{\text{J}}(t)=(0,0)$ for any $t\geq 0$.
@@
 $\mathbb{E}\left[T^{(0,0)}_{(1,0)}\right]=\frac{1}{\lambda}\left(\frac{1}{(1-\rho)^{2}}-1\right)=\frac{1}{\lambda}\frac{\rho(2-\rho)}{(1-\rho)^{2}}.$ (3.12)
-Next, consider a two-queue JSQ system $\bm{q}^{\text{J}}(t)=(q_{1}(t),q_{2}(t))$ (with an arrival rate $\lambda$ and each queue has a service rate $\mu$) and the product chain $\bm{q}^{\text{M}}(t)=(q^{\text{M}}_{1}(t),q^{\text{M}}_{2}(t))$ of two independent $M/M/1$ queues defined previously, started from the same initial state $\bm{q}^{\text{J}}(0)=\bm{q}^{\text{M}}(0)=(1,0)$.
-By Proposition 4, there exists a copuling such that
+Next, consider a two-queue JSQ system $\bm{q}^{\text{J}}(t)=(q^{\text{J}}_{1}(t),q^{\text{J}}_{2}(t))$ (with an arrival rate $\lambda$ and each queue has a service rate $\mu$) and the product chain $\bm{q}^{\text{M}}(t)=(q^{\text{M}}_{1}(t),q^{\text{M}}_{2}(t))$ of two independent $M/M/1$ queues defined previously, started from the same initial state $\bm{q}^{\text{J}}(0)=\bm{q}^{\text{M}}(0)=(1,0)$.
+By Proposition 4, there exists a coupling such that
 $T_{0}=\inf\left\{t\geq 0:\bm{q}^{\text{J}}(t)=(0,0)\right\}\leq\inf\left\{t\geq 0:\bm{q}^{\text{M}}(t)=(0,0)\right\}=T^{(0,0)}_{(1,0)}\qquad\text{almost surely.}$